# SD2: Self-Distilled Sparse Drafters

### v1

@@ 3.2 MACs analysis
While SD2 drafters perform comparably to dense models, it remains unclear whether they offer practical benefits. Unstructured sparsity is not well suited for GPUs and does not typically yield a tangible latency decrease. Nevertheless, future hardware and algorithms may offer improved unstructured sparsity support. From this perspective, an analysis based on MACs provides a theoretical limit of the benefits one can expect from unstructured sparsity.
Figure 5: MAL vs. MACs for layer-pruned and sparse draft models. Particularly notable are the Qwen-2.5 unstructured sparse drafters which approach iso-MAC performance compared to the dense models.
In Figure 5, we plot the MAL of our models trained with SD2 versus their corresponding MACs. The unstructured sparse Qwen-2.5 drafters nearly achieve iso-MAC performance improvements compared to their dense counterparts. However, additional MACs associated with the lm-head layer prevent the unstructured sparse models from directly outperforming the dense drafters in terms of MAL vs. MACs. In Figure 1 we demonstrate that our best performing SD2 draft models provide the highest theoretical improvement factor of all draft model candidates examined.
@@ 3.3 Compressed sparse drafters
While our MACs analysis highlights the potential theoretical benefits of SD2, unstructured sparse models remain unpractical presently. As noted in Equation 3, a crucial property to consider when evaluating speculative decoding draft models is $c$, the ratio of the draft to target model latency. To realize practical benefits of SD2 on commodity hardware available today, we require efficient sparse representations and kernels which offer lower latency than the dense baselines. To this end, we benchmark the following compression schemes:
  • 2:4 FP16 – 2:4 sparsity with FP16 weights and activations;
  • 2:4 W8A8 – 2:4 sparsity with Int8 weights and activations;
  • Marlin – Sparse-Marlin with Int4 weights and FP16 activations, including the lm-head layer;
  • Marlin FP16 head – Sparse-Marlin with Int4 weights and FP16 activations, excluding the lm-head layer.
2:4 sparsity was introduced by  Mishra et al. (2021) and is compatible with NVIDIA GPUs. Sparse-Marlin (Frantar et al., 2024) is a state-of-the-art kernel that supports 2:4 sparsity combined with 4-bit quantization. See Appendix C for details on our our quantization and benchmarking methodology. We compare the improvement factor, MAL and latencies of the various compression schemes in Figure 7. Despite lower latencies for our compressed drafters, the dense model baselines offer a better end-to-end improvement. Sparse-Marlin with FP16 head offers the most competitive improvement factor amongst our compressed drafters.
@@ 4 Related work
Speculative decoding with greedy verification was initially proposed by Stern et al. (2018). Speculative sampling was introduced concurrently by Leviathan et al. (2023) and Chen et al. (2023). Since these initial investigations, several works have proposed modifications and refinements to the original framework. In Medusa (Cai et al., 2024), the authors reformulated used a single model with multiple lm-head layers, each of which predicts multiple draft token candidates across each draft token position, introducing the first instance of draft token-trees. Specifically related to our work, Medusa also used self-distillation to obtain a fine-tuning dataset. Chen et al. (2024a) examined speculative decoding for long context lengths, finding that constant KV-cache sizes lead to improved acceleration. EAGLE (Li et al., 2024c, b) proposed an efficient draft model consisting of a single layer auto-regression head trained from scratch and the frozen embedding and lm-head layers from the target model. Token trees were further investigated by Miao et al. (2024) in which a parallel decoding algorithm was also proposed. Sun et al. (2024a) introduced a hierarchical framework in which efficient KV-cache implementations are used for drafting. Speculative decoding was combined with early-exit mechanisms for drafting by Zhang et al. (2024a) and Liu et al. (2024a). We note that optimizations such as draft token-trees and KV-cache compression are orthogonal to SD2 and future work could consider their integration with our method.
@@
The surprising result that a significant fraction of decoder blocks can be pruned from LLMs without incurring excessive quality reduction was examined in detail in several works (Jha et al., 2024; Sun et al., 2024b; Yang et al., 2024a; Men et al., 2024). In particular, these works found that the “middle” decoder blocks are the most amenable to pruning. However, abstract reasoning tasks were noted to be disproportionality affected by layer-pruning. We specifically highlight Gromov et al. (2024) which introduced the angular cosine distance metric that is used for our layer-pruned baselines.
@@
Fine-grained sparsity for LLMs has received significant attention from the research community. Efficient one-shot pruning methods such as SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2023), and Plug-and-Play (Zhang et al., 2023) proposed re-framing the pruning procedure as a layer-wise reconstruction of the original model weights, using the magnitudes and/or activations to determine the saliency of individual weights. Fang et al. (2024) proposed freezing the model weights and training the mask alone to efficiently obtain 2:4 sparse LLMs. OWL (Yin et al., 2023b) and AlphaPruning (Lu et al., 2024) proposed non-uniform layer-wise sparsity distributions which outperform uniform distributions in terms of perplexity and downstream task accuracy. Despite these efforts, sparse LLMs still fall short of their dense counterparts in many respects (Jaiswal et al., 2023; Thangarasa et al., 2023a; Yin et al., 2023a).
@@ 5 Conclusion
We introduce the SD2 methodology for obtaining fine-grained sparse draft models. SD2 consists of self-data distillation, one-shot pruning, and sparse fine-tuning. We find that self-data distillation effectively aligns draft models, even if their target is from a different model family such as in the UAG setting. While SD2 shows significant theoretical benefits based MACs, currently available optimized kernels and compressed representations fail to offer practical end-to-end latency benefits. See Section 6 for a discussion of limitations and future work. We hope this work will inspire further efforts to develop low-latency sparse kernels and encourage software/hardware co-design for efficient LLM inference.
6 Limitations and future work
The effectiveness of SD2 in improving inference efficiency depends on software and hardware optimized for sparsity. Sparse-Marlin is designed for Ampere-series NVIDIA GPUs, and evaluating its performance on alternative hardware, such as TPUs or custom accelerators, is crucial for broader adoption. While SD2 requires significant compute for fine-tuning and self-data distillation, this investment could further enhance the quality of dense models, particularly in speculative decoding. Several promising research directions emerge from this work. One is integrating quantization-aware training (Chen et al., 2024b), draft-token trees, and compressed KV-cache implementations (Shi et al., 2024) with SD2 to improve memory and compute efficiency. Another is benchmarking SD2 on hardware optimized for unstructured sparsity, such as Cerebras’ wafer-scale engine. Incorporating advanced fine-tuning strategies like speculative knowledge distillation (Xu et al., 2024a) or square-head distillation (Kurtic et al., 2023) could further draft model quality. Finally, combining structured pruning with fine-grained sparsity and quantization presents an opportunity to reduce inference-time latency while maintaining accuracy, making sparsity-aware methods like SD2 more widely applicable across diverse hardware architectures.
Impact Statement
@@ 6 Limitations and future work
Appendix A Hyperparameter settings and implementation details
As noted above, the core components of SD2 consist of: one-shot pruning, self-data distillation, and sparse fine-tuning.
A.1 One-shot pruning
@@
We prune our draft model candidates to 0.5, 0.66, and 0.75 unstructured sparsity in addition to 2:4 sparsity. For our experiments with non-uniform layer-wise sparsity distributions, we explore both OWL and our proposed angular distance layer-wise distribution, the results of which are presented in Appendix B.
However, despite apparent advantages in terms of perplexity when measured on WikiText-V2 (Merity et al., 2016), we find that the non-uniform layer-wise sparsity distributions considered offer little benefit to downstream evaluation tasks or MAL during speculative decoding. Furthermore, since non-uniform layer-wise sparsity distribution pose a challenge for inference systems that leverage pipelining with continuous batching, we opt to use the more straightforward uniform layer-wise sparsity distribution for all results included in Section 3.
A.2 Self-data distillation
Our self-data distillation follows the original implementation [^5] https://github.com/sail-sg/sdft of  Yang et al. (2024b). We generate the distilled labels using 16 bit float precision for the model weights and a maximum generation length of 4096 tokens. We sample from the target model with top-p sampling with $p=1.0$ and a temperature of 0.9. Using Llama-3.1-70B, we use self-data distillation to produce aligned fine-tuning datasets for Opus-100 translation (Tiedemann and Thottingal, 2020) [^6] https://huggingface.co/datasets/Helsinki-NLP/opus-100 and MathInstruct-V2 (Toshniwal et al., 2024) datasets. Using Qwen-2.5-72B-Instruct, we distil MathInstruct-V2.
@@ A.2 Self-data distillation
In addition to these datasets, we leverage publicly available synthetic datasets produced using the self-synthesis method Magpie (Xu et al., 2024b) for both Llama-3.1-70B-Instruct [^7] https://huggingface.co/datasets/Magpie-Align/Magpie-Llama-3.1-Pro-MT-300K-Filtered and Qwen-2.5-72B-Instruct [^8] https://huggingface.co/datasets/Magpie-Align/Magpie-Qwen2.5-Pro-300K-Filtered target models. As the Magpie datasets are synthetically generated by the target models, further alignment with self-data distillation is redundant.
A.3 Sparse fine-tuning
@@
We optimize our models using the default AdamW optimizer (Loshchilov and Hutter, 2017) with 1.0e-08, 0.9, and 0.999 for the $\varepsilon$, $\beta_{1}$, and $\beta_{2}$ hyperparameters, respectively. For most of our models, we train them with 16,000 optimizer steps with a batch size of 8 for a total of 128,000 samples. The one exception to the above are the 75% sparse models, which we find benefit from an extended training duration of 32,000 optimizer steps with a batch size of 8 for a total of 256,000 samples.
@@
Appendix B Non-uniform layer-wise sparsity distributions
This section highlights our analysis of non-uniform layer-wise sparsity distributions such as OWL (Yin et al., 2023b) and a novel distribution inspired by the angular cosine distance measure from  Gromov et al. (2024). Ultimately, we found that the non-uniform layer-wise sparsity distributions examined resulted in improved perplexity on WikiText-V2 and a modest improvement on the OpenLLM Leaderboard V1 suite of benchmarks; however, we did not find a statistically significant improvement to MAL during speculative decoding.
OWL allocates sparsity to layers proportional to their outlier ratio. The outlier ratio was inspired by the observation that the activations of LLMs often contain large outliers (Dettmers et al., 2022). Specifically, for a weight matrix $\mathbf{W}$ of shape ($C_{out}$, $C_{in}$) and input $X$, OWL defines the outlier score of $\mathbf{W}_{i,j}$ as $A_{i,j}=\left\|X_{j}\right\|\cdot\left|\mathbf{W}_{i,j}\right|$ across all tokens in a calibration dataset. The sparsity ratio is defined as $S^{l}\propto 1-D^{l}\quad S^{l}\in\{S-\lambda,S+\lambda\}$ where $S\in[0,1]$ is the uniform sparsity target, $\lambda$ is a hyperparameter which constraints the maximum and minimum layer sparsities, and $D^{l}$ is the outlier distribution for a single layer calculated as follows:
$D^{l}=\frac{\sum_{i=1}^{C_{out}}\sum_{j=1}^{C_{in}}\mathbf{1}_{\zeta}(A^{l}_{i,j})}{C_{in}C_{out}}$ (8)
In the above, $\bar{A}^{l}$ is the mean of $A^{l}$, $M$ is a hyperparameter which defines the magnitude of outliers compared to the average activation, and $\mathbf{1}_{\zeta}$ is the indicator function defined as:
$\mathbf{1}_{\zeta}=\begin{cases}1&A_{i,j}>M\cdot\bar{A}^{l}\\ 0&\text{otherwise}\end{cases}$ (9)
@@ Appendix B Non-uniform layer-wise sparsity distributions
Formally, for a decoder with $n$ layers, target global decoder sparsity $S\in[0,1]$, and angular distances $\mathbf{D}\in[D^{1},...,D^{N}]$ where $D^{i}$ is defined by Equation 6 with $n=1$, we allocate the sparsity of each block as:
$S^{i}=1-\left((1-S)*\frac{D^{i}}{\sum_{j=1}^{n}\left|D^{j}\right|}\right)$ (10)
@@
Figure 6: Comparison between non-uniform layer-wise sparsity distributions OWL and our proposed angular distance distribution. In Figures 6, 6 and 6, we report the WikiText-V2 Perplexity for Llama-3.2 1B and 3B and Llama-3.1 8B at 50, 60, 70, 80, and 90% sparsity. At high sparsities, the angular distance distribution outperforms OWL, particularly for the 1B model. In Figures 6, 6 and 6, we report the ARC-C accuracy for the same models and sparsities. Despite outperforming OWL in terms of perplexity, we find that OWL achieves slightly higher scores on this benchmark. Further, neither distribution yields statistically significant improvements to AR in our limited experiments.
Appendix C Compressed sparse drafter quantization and benchmarking
For quantization, we follow a simple Post-Training Quantization (PTQ) scheme. For Int8 weights and activations combined with 2:4 sparsity (W8A8), we use dynamic per-token quantization for the activations and static symmetric per-row quantization for the weights. For Sparse-Marlin, we use static symmetric per-group quantization with a group size of 128. We use the open-source kernels available from torchao (torchao maintainers and contributors, 2024) to benchmark our compressed draft models. Benchmarking was completed on a 4 $\times$ NVIDIA A6000 GPU node. We compile each model using torch.compile with mode, dynamic, and full-graph arguments set to max-autotune, True, and True, respectively. We use the default dynamic KV-cache implementation from HuggingFace’s transformers library as the static cache implementation remains incomplete at the time of writing.
Latencies were recorded with the torch.utils.benchmark tools by repeatedly measuring the latency of the model’s forward pass until at least 20 seconds had elapsed. We record the median latency measured across all forward passes completed during this duration for both the draft models and target model. From the median latencies, we obtain $c$, the cost factor required to calculate Equation 3. We report the improvement factor for the various compression schemes in Figure 7(a).
Notably, we did not fine-tune the quantized models after quantization using a Quantization-Aware Training (QAT) workflow. We believe that fine-tuning with QAT or using one of the various more sophisticated quantization schemes would further boost the quality of these models.
(a) Practical improvement factor
(b) Mean Accepted Length (MAL)
(c) Latency
Figure 7: Compressed SD2 draft model properties. In Figure 7(a), we find that despite latency improvements for our compressed draft models, the higher MAL of the dense draft models results in a better overall improvement factor. In Figure 7(b), we find that our PTQ quantization process significantly reduces the MAL of our compressed models; however, in Figure 7(c) this quantization leads to significant reductions in latency versus the dense baseline. Notably, the 2:4 FP16 kernel maintains MAL but increases latency compared to the dense draft models, likely due to metadata overhead at batch size 1. We record latencies using a batch size and sequence length of 1.
Appendix D Detailed SpecBench results

### v2

@@ 3.2 MACs analysis
The prior section demonstrate that SD2 drafters have comparable quality to dense models, it remains unclear whether their efficiency gains yield offer practical benefits. Unstructured sparsity is not well suited for GPUs and tangible latency improvements require specialized kernels. An analysis based on MACs provides a theoretical limit of the benefits one can expect from unstructured sparsity.
In Figure 2(b), we plot the MAL of our models trained with SD2 versus their corresponding MACs. The unstructured sparse Qwen-2.5 drafters nearly achieve iso-MAC performance improvements compared to their dense counterparts. However, additional MACs associated with the LM head prevent the unstructured sparse models from directly outperforming the dense drafters in terms of MAL vs. MACs. In Figure 5 we demonstrate that our best performing SD2 draft models provide the highest theoretical improvement factor of all draft model candidates examined.
@@ 3.3 Compressed sparse drafters
As noted in Equation 3, a crucial property to consider when evaluating speculative decoding draft models is $c$, the ratio of the draft to target model latency. To realize practical benefits of SD2 on commodity hardware available today, we require efficient sparse representations and kernels which offer lower latency than the dense baselines. To this end, we benchmark our SD2 unstructured draft models using nm-vLLM (Neural Magic, 2024; Kwon et al., 2023) and compare them with dense and layer-pruned draft models in Figure 1. SD2 drafters outperform layer-pruned models and even dense drafters in the 3B model size category. While the smallest dense model in each model family achieves a higher overall improvement factor, these results highlight the potential for sparsity to enable the use of larger draft models. As foundation model sizes continue to grow, increasing the capacity of draft models may be necessary to maintain sufficient MAL and support novel use-cases such as scaling test-time compute. See Appendix E for details on our benchmarking methodology.
@@ 4 Related work
Speculative decoding with greedy verification was initially proposed by Stern et al. (2018). Speculative sampling was introduced concurrently by Leviathan et al. (2023) and Chen et al. (2023). Since these initial investigations, several works have proposed modifications and refinements to the original framework. In Medusa (Cai et al., 2024), the authors reformulated used a single model with multiple lm-head layers, each of which predicts multiple draft token candidates across each draft token position, introducing the first instance of draft token-trees. Specifically related to our work, Medusa also used self-distillation to obtain a fine-tuning dataset. Sadhukhan et al. (2024) examined speculative decoding for long context lengths, finding that constant KV-cache sizes lead to improved acceleration. EAGLE (Li et al., 2024c, b) proposed an efficient draft model consisting of a single layer auto-regression head trained from scratch and the frozen embedding and lm-head layers from the target model. Token trees were further investigated by Miao et al. (2024) in which a parallel decoding algorithm was also proposed. Sun et al. (2024a) introduced a hierarchical framework in which efficient KV-cache implementations are used for drafting. Speculative decoding was combined with early-exit mechanisms for drafting by Zhang et al. (2024a) and Liu et al. (2024a). We note that optimizations such as draft token-trees and KV-cache compression are orthogonal to SD2 and future work could consider their integration with our method.
@@
The surprising result that a significant fraction of decoder blocks can be pruned from LLMs without incurring excessive quality reduction was examined in detail in several works (Jha et al., 2024; Sun et al., 2025; Yang et al., 2024a; Men et al., 2024). In particular, these works found that the “middle” decoder blocks are the most amenable to pruning. However, abstract reasoning tasks were noted to be disproportionality affected by layer-pruning. We specifically highlight Gromov et al. (2025) which introduced the angular cosine distance metric that is used for our layer-pruned baselines.
@@
Fine-grained sparsity for LLMs has received significant attention from the research community. Efficient one-shot pruning methods such as SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024b), and Plug-and-Play (Zhang et al., 2024b) proposed re-framing the pruning procedure as a layer-wise reconstruction of the original model weights, using the magnitudes and/or activations to determine the saliency of individual weights. Fang et al. (2024) proposed freezing the model weights and training the mask alone to efficiently obtain 2:4 sparse LLMs. OWL (Yin et al., 2024b) and AlphaPruning (Lu et al., 2024) proposed non-uniform layer-wise sparsity distributions which outperform uniform distributions in terms of perplexity and downstream task accuracy. Despite these efforts, sparse LLMs still fall short of their dense counterparts in many respects (Jaiswal et al., 2024; Thangarasa et al., 2023; Yin et al., 2024a).
@@ 5 Conclusion
We introduce the SD2 methodology for obtaining fine-grained sparse draft models. SD2 consists of self-data distillation, one-shot pruning, and sparse fine-tuning. We find that self-data distillation effectively aligns draft models, even if their target is from a different model family such as in the UAG setting. SD2 shows significant theoretical benefits based MACs. When paired with efficient sparse kernels, SD2 outperforms layer-pruned models at all model sizes and compression ratios evaluated. For the 3B draft models, SD2 improves end-to-end speculative decoding latency compared to dense draft models, revealing the potential for using larger drafters. See Appendix B for a discussion of limitations and future work. We hope this work will inspire further efforts to develop low-latency sparse kernels and encourage software/hardware co-design for efficient LLM inference.
6 Acknowledgements
We would like to acknowledge the helpful feedback of Ganesh Venkatesh during our initial explorations that led to this work. ML and YI gratefully acknowledge the support of Alberta Innovates (ALLRP-577350-22, ALLRP-222301502), the Natural Sciences and Engineering Research Council of Canada (NSERC) (RGPIN-2022-03120, DGECR-2022-00358), and Defence Research and Development Canada (DGDND-2022-03120). YI is supported by a Schulich Research Chair.
Appendix A Impact Statement
@@
Appendix B Limitations and future work
The effectiveness of SD2 in improving inference efficiency depends on software and hardware optimized for sparsity. The nm-vLLM unstructured sparse kernel we use to benchmark is optimized for Ampere-series NVIDIA GPUs, and evaluating its performance on alternative hardware, such as TPUs or custom accelerators, is crucial for broader adoption. While SD2 requires significant compute for fine-tuning and self-data distillation, this investment could further enhance the quality of dense models, particularly in speculative decoding. Several promising research directions emerge from this work. One is integrating quantization-aware training (Chen et al., 2025), draft-token trees, and compressed KV-cache implementations (Luohe et al., 2024) with SD2 to improve memory and compute efficiency. Another is benchmarking SD2 on hardware optimized for unstructured sparsity, such as Cerebras’ wafer-scale engine. Incorporating advanced fine-tuning strategies like speculative knowledge distillation (Xu et al., 2024a) or square-head distillation (Kurtic et al., 2023) could further improve sparse draft model quality. Finally, combining structured pruning with fine-grained sparsity and quantization presents an opportunity to reduce inference-time latency while maintaining accuracy, making sparsity-aware methods like SD2 more widely applicable across diverse hardware architectures.
Appendix C Hyperparameter settings and implementation details
The core components of SD2 consist of: one-shot pruning, self-data distillation, and sparse fine-tuning.
C.1 One-shot pruning
@@
We prune our draft model candidates to 0.5, 0.66, and 0.75 unstructured sparsity in addition to 2:4 sparsity. For our experiments with non-uniform layer-wise sparsity distributions, we explore both OWL and our proposed angular distance layer-wise distribution, the results of which are presented in Appendix D.
However, despite apparent advantages in terms of perplexity when measured on WikiText-V2 (Merity et al., 2017), we find that the non-uniform layer-wise sparsity distributions considered offer little benefit to downstream evaluation tasks or MAL during speculative decoding. Furthermore, since non-uniform layer-wise sparsity distribution pose a challenge for inference systems that leverage pipelining with continuous batching, we opt to use the more straightforward uniform layer-wise sparsity distribution for all results included in Section 3.
C.2 Self-data distillation
Our self-data distillation follows the original implementation [^5] https://github.com/sail-sg/sdft of  Yang et al. (2024b). We generate the distilled labels using 16 bit float precision for the model weights and a maximum generation length of 4096 tokens. We sample from the target model with top-p sampling with $p=1.0$ and a temperature of 0.9. Using Llama-3.1-70B, we use self-data distillation to produce aligned fine-tuning datasets for Opus-100 translation (Tiedemann and Thottingal, 2020) [^6] https://huggingface.co/datasets/Helsinki-NLP/opus-100 and MathInstruct-V2 (Toshniwal et al., 2025) datasets. Using Qwen-2.5-72B-Instruct, we distil MathInstruct-V2.
@@ C.2 Self-data distillation
In addition to these datasets, we leverage publicly available datasets produced using the self-synthesis method Magpie (Xu et al., 2024b) for both Llama-3.1-70B-Instruct [^7] https://huggingface.co/datasets/Magpie-Align/Magpie-Llama-3.1-Pro-MT-300K-Filtered and Qwen-2.5-72B-Instruct [^8] https://huggingface.co/datasets/Magpie-Align/Magpie-Qwen2.5-Pro-300K-Filtered target models. As the Magpie datasets are generated by the target models, further alignment with self-data distillation is redundant.
C.3 Sparse fine-tuning
@@
We optimize our models using the default AdamW optimizer (Loshchilov and Hutter, 2018) with 1.0e-08, 0.9, and 0.999 for the $\varepsilon$, $\beta_{1}$, and $\beta_{2}$ hyperparameters, respectively. For most of our models, we train them with 16,000 optimizer steps with a batch size of 8 for a total of 128,000 samples. The one exception to the above are the 75% sparse models, which we find benefit from an extended training duration of 32,000 optimizer steps with a batch size of 8 for a total of 256,000 samples.
@@
C.4 Computational Resources
We use an internal cluster to run our experiments on a node with four Nvidia H100-80GB-HBM3 GPUs, 192 Intel Xeon Platinum 8468 CPUs, and 2TB of RAM. Pruning was completed using a single GPU in less than 30 minutes per model. Our fine-tuning datasets were generated in approximately 6-8 hours per dataset. For sparse fine-tuning, we use a single GPU with the following approximate runtimes: 10 and 21 hours for 1B and 3B, respectively, for the 50% and 66% sparse models. The 75% sparse model fine-tuning took approximately 20 and 42 hours for 1B and 3B, respectively, due to the increased number of optimization steps.
Appendix D Non-uniform layer-wise sparsity distributions
This section highlights our analysis of non-uniform layer-wise sparsity distributions such as OWL (Yin et al., 2024b) and a novel distribution inspired by the angular cosine distance measure from  Gromov et al. (2025). Ultimately, we found that the non-uniform layer-wise sparsity distributions examined resulted in improved perplexity on WikiText-V2 and a modest improvement on the OpenLLM Leaderboard V1 suite of benchmarks; however, we did not find a statistically significant improvement to MAL during speculative decoding.
OWL allocates sparsity to layers proportional to their outlier ratio. The outlier ratio was inspired by the observation that the activations of LLMs often contain large outliers (Dettmers et al., 2022). Specifically, for a weight matrix $\mathbf{W}$ of shape ($C_{out}$, $C_{in}$) and input $X$, OWL defines the outlier score of $\mathbf{W}_{i,j}$ as $A_{i,j}=\left\|X_{j}\right\|\cdot\left|\mathbf{W}_{i,j}\right|$ across all tokens in a calibration dataset. The sparsity ratio is defined as $S^{l}\propto 1-D^{l}\quad S^{l}\in\{S-\lambda,S+\lambda\}$ where $S\in[0,1]$ is the uniform sparsity target, $\lambda$ is a hyperparameter which constraints the maximum and minimum layer sparsities, and $D^{l}$ is the outlier distribution for a single layer calculated as
$D^{l}=\frac{\sum_{i=1}^{C_{out}}\sum_{j=1}^{C_{in}}\mathbf{1}_{\zeta}(A^{l}_{i,j})}{C_{in}C_{out}}.$ (8)
In the above, $\bar{A}^{l}$ is the mean of $A^{l}$, $M$ is a hyperparameter which defines the magnitude of outliers compared to the average activation, and $\mathbf{1}_{\zeta}$ is the indicator function defined as
$\mathbf{1}_{\zeta}=\begin{cases}1&A_{i,j}>M\cdot\bar{A}^{l}\\ 0&\text{otherwise}.\end{cases}$ (9)
@@ Appendix D Non-uniform layer-wise sparsity distributions
Formally, for a decoder with $n$ layers, target global decoder sparsity $S\in[0,1]$, and angular distances $\mathbf{D}\in[D^{1},...,D^{N}]$ where $D^{i}$ is defined by Equation 6 with $n=1$, we allocate the sparsity of each block as
$S^{i}=1-\left((1-S)*\frac{D^{i}}{\sum_{j=1}^{n}\left|D^{j}\right|}\right).$ (10)
@@
Figure 6: Comparison between non-uniform layer-wise sparsity distributions OWL and our proposed angular distance distribution. In Figures 6, 6 and 6, we report the WikiText-V2 Perplexity for Llama-3.2 1B and 3B and Llama-3.1 8B at 50, 60, 70, 80, and 90% sparsity. At high sparsities, the angular distance distribution outperforms OWL, particularly for the 1B model. In Figures 6, 6 and 6, we report the ARC-C accuracy for the same models and sparsities. Despite outperforming OWL in terms of perplexity, we find that OWL achieves slightly higher scores on this benchmark. Further, neither distribution yields statistically significant improvements to acceptance rate in our limited experiments.
Appendix E Benchmarking details
Benchmarking was completed on a 4 $\times$ NVIDIA A6000 GPU node with 88 Intel Xeon E5-2699 v4 CPUs using nm-vLLM (Neural Magic, 2024) version 0.5.0. The throughput of our draft and target models is measured using a single request of 300 input tokens and 256 output tokens, the average token counts from the SpecBench benchmark. We repeat the request 10 times to minimize variance in the measurement. The cost factor, $c$, is calculated as the ratio of the draft and target model tokens per second including both input and output tokens. Using MAL from our SpecBench results, we derive the improvement factor per Equation 3. We report the improvement factor and throughput for dense, layer-pruned, and SD2 unstructured drafters in Figures 1 and 7, respectively. The SD2 1.5B and 3B drafters outperform both dense and layer-pruned models in terms of end-to-end latency improvements.
In Figure 8 we report the improvement factor for 10 concurrent requests following a similar benchmarking approach as per the above. Notably, SD2 still provides the highest improvement factor for the 3B draft model category, but the 1.5B dense drafter slightly outperforms SD2 in this setting.
Figure 7: Throughput (tok/s) vs. sparsity by model and variant. Layer-pruned draft models offer ultra-low latency; however, their reduced MAL results in an overall lower improvement factor compared to SD2 unstructured sparse models. All models are benchmarked using nm-vLLM. Throughput values include both input and output tokens.
Figure 8: Improvement factor of dense, layer-pruned, and SD2 unstructured draft models when benchmarked with 10 concurrent requests. Llama and Qwen models drafting for Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct, respectively. Even in this setting, SD2 drafters outperform layer-pruned draft models and dense drafters in the 3B model size category.
Appendix F Detailed SpecBench results